# The double spherical cap rearrangement of planar sets

### v1

@@ Abstract.
This paper is devoted to the proof of an isoperimetric property of the double spherical cap rearrangement of planar sets under the assumption of disconnection of non-trivial spherical slices. Additionally, the higher-dimensional case is briefly discussed; in particular, an explicit counterexample is given, thus explaining why an analogous result cannot hold.
@@ 1. Introduction
A crucial tool in the theory of shape optimization is given by the rearrangements of sets, which allow us to keep some properties of the original set, while improving other aspects. A quite complete guide about this topic is surely contained in [10] and in the references therein. In particular, when symmetrizing a set in some way, it is often useful to know that the rearrangement has the same volume of the original set, while decreasing the perimeter; indeed, it is not rare the case where, in solving some variational problem, one would like to reduce the searching field to sets that have some symmetry property. An iconical example is the isoperimetric problem, where a key step in the proof by De Giorgi has been the proof of a rigidity property of Steiner’s inequality (see, for example [6]). Lately, a very productive research field is developping around quantitative isoperimetric inequalities (see for instance [7, 5, 3, 9] and the references therein). In this setting, a very useful symmetrization is given by the spherical rearrangement (first introduced in [12]), which allows one to decrease the perimeter, while keeping the volume of the intersection between the set and all balls centered at the origin. This last symmetrization is widely studied in [4], where the authors address in particular the question of rigidity. However, in some situations the spherical cap rearrangement is not useful. Indeed, this symmetrization does not keep the barycenter and this could be a problem. This is the case, for instance, of problems involving the barycentric asymmetry, that is, the volume of the symmetric difference between a set and the ball with the same volume and the same barycenter (see for instance [7, 8, 3, 9]). The double spherical cap rearrangement is a different symmetrization that can be helpful in such situations. Roughly speaking, while in the spherical rearrangement of a set $E$ all the (N1)(N-1)-dimensional area of $E\cap\partial B(r)$ is moved towards a fixed direction, say $e_{N}$, in the double spherical cap rearrangement this area is split in two equal parts, one of which is moved towards the direction $e_{N}$, while the other one is moved towards the direction $e_{N}$. The advantage of this rearrangement is that not only it keeps the volume of $E$, but also its barycenter, as soon as the procedure is centered at the barycenter of the set in question; the disadvantage, however, is that the perimeter is not decreased in general. In this paper we will present this symmetrization and study its properties, showing in particular that the perimeter decreases in dimension $2$ if all non-trivial slices are disconnected.
@@ 2. Preliminary notions
By [4, Lemma 3.2], in dimension $n=2$ the following expressions hold for every $x\in\mathbb{R}^{2}_{0}$:
@@
By [4, Lemma 3.4], for every $\varphi\in C^{1}_{c}(\mathbb{R}^{2}_{0},\mathbb{R}^{2})$ the following hold:
@@
We close this section of preliminary results with a version of a result by Vol’pert (see [4, Theorem 3.7]):
@@ 3. Technical results
We collect in this section some technical results that will be the core of the proof of Theorem 1.3. The first one is a particular case of [4, Lemma 4.1]; for the sake of completeness, however, we show here the proof for the case $n=2$.

### v2

@@ Abstract.
In the theory of shape optimization the rearrangements of sets are a key concept, because they allow us to keep some properties of the original set while improving other aspects. This paper is devoted to the proof of an isoperimetric property of the double spherical cap rearrangement of planar sets. In particular, we prove that, under the assumption of disconnection of non-trivial spherical slices, the rearranged set has a lower perimeter than the original one. In the general case, the symmetrized set does not decrease the perimeter, but we show that the “excess” is bounded above by $2\mathcal{H}^{1}(\Gamma)$, where $\Gamma$ denotes the set of radii such that the spherical slice is a non-trivial arc of circle. Additionally, the higher-dimensional case is briefly discussed; in particular, an explicit counterexample is given, thus explaining why an analogous result cannot hold. The main reason for this is that, in dimension $N=3$ or higher, the union of two spherical caps of equal size does not minimize the (N2)(N-2)-dimensional measure of the boundary.
Key words and phrases:
spherical rearrangements, isoperimetric inequality, rigidity, symmetrization
2020 Mathematics Subject Classification:
28A75
@@ 1. Introduction
A crucial tool in the theory of shape optimization is given by the rearrangements of sets, which allow us to keep some properties of the original set while improving other aspects. A quite complete guide about this topic is surely contained in [12] and in the references therein. In particular, when symmetrizing a set in some way, it is often useful to know that the rearrangement has the same volume of the original set, while decreasing the perimeter; indeed, it is not rare the case where, in solving some variational problem, one would like to reduce the searching field to sets that have some symmetry property. An iconical example is the isoperimetric problem, where a key step in the proof by De Giorgi has been the proof of a rigidity property of Steiner’s inequality (see, for example [8]). Lately, a very productive research field is developping around quantitative isoperimetric inequalities (see for instance [9, 7, 3, 11] and the references therein). In this setting, a very useful symmetrization is given by the spherical rearrangement (first introduced in [14]), which allows one to decrease the perimeter, while keeping the volume of the intersection between the set and all balls centered at the origin. This last symmetrization is widely studied in [6], where the authors address in particular the question of rigidity. However, in some situations the spherical cap rearrangement is not useful. Indeed, this symmetrization does not keep the barycenter and this could be a problem. This is the case, for instance, of problems involving the barycentric asymmetry, that is, the volume of the symmetric difference between a set and the ball with the same volume and the same barycenter (see for instance [9, 10, 3, 11]). The double spherical cap rearrangement is a different symmetrization that can be helpful in such situations. Roughly speaking, while in the spherical rearrangement of a set $E$ all the (N1)(N-1)-dimensional area of $E\cap\partial B(r)$ is moved towards a fixed direction, say $e_{N}$, in the double spherical cap rearrangement this area is split in two equal parts, one of which is moved towards the direction $e_{N}$, while the other one is moved towards the direction $e_{N}$. The advantage of this rearrangement is that not only it keeps the volume of $E$, but also its barycenter, as soon as the procedure is centered at the barycenter of the set in question; the disadvantage, however, is that the perimeter is not decreased in general. In this paper we will present this symmetrization and study its properties, showing in particular that the perimeter decreases in dimension $2$ if all non-trivial slices are disconnected.
Remark 1.1.
This kind of symmetrization was first introduced in [4], where Bonnesen works with convex sets. In particular, he proves that, by centering this rearrangement in the right point, the symmetrized set has a lower perimeter than the original set. This was used in [5], where the author refers to this construction as the “Bonnesen semicircular symmetrization”, summarizing very well its main properties. In a way, this paper generalizes what Bonnesen proved. Indeed, the “special point”, where the symmetrization is centered, is the center of the annulus of minimal witdh containing the boundary of the set, which is actually such that all the non-trivial spherical slices are disconnected.
@@
Acknowledgments.
The author would like to thank Giorgio Saracco, for pointing out references [4, 5]. The author is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM).
@@ 2. Preliminary notions
By [6, Lemma 3.2], in dimension $n=2$ the following expressions hold for every $x\in\mathbb{R}^{2}_{0}$:
@@
By [6, Lemma 3.4], for every $\varphi\in C^{1}_{c}(\mathbb{R}^{2}_{0},\mathbb{R}^{2})$ the following hold:
@@
We close this section of preliminary results with a version of a result by Vol’pert (see [6, Theorem 3.7]):
@@ 3. Technical results
We collect in this section some technical results that will be the core of the proof of Theorem 1.4. The first one is a particular case of [6, Lemma 4.1]; for the sake of completeness, however, we show here the proof for the case $n=2$.